OpenStudy (anonymous):

use the formula to evaluate the series: 3+6+12+24+48+96

OpenStudy (anonymous):

this is a GP with a = 3 and r = 2 sum of 6 terms = a (r^6 - 1) / r-1 = 3 (2^6 -1) / 2-1 = 3*63 / 1 =189